Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Paoletti, Tringali, Fiori - 18:00 Wednesday 07 November 2018 (43450) Print this report
Test of static mirror charge at NE and WE

We injected a sinusoidal signal to each one of the four coils (DR, UR, DL, UL) of the NE and then WE payload.  We used a 1Vrms at 15Hz set at the generator's output (but ~10% less applied to the coils due to a 9/10 partition for monitoring purpose).

Note that the Coils are used with the two wires shot-circuited, and the voltage sine is applied between these two wires and the Virgo ground (tower's vacuum chamber). When we apply the voltage to one coil the other three are connected.

The table below summarizes the measurements. The behaviour of both mirrors is consistent with that of a charged mirror. To be better investigated.

We noticed that when injecting at the WE, the noise floor in the region 20 to 40Hz sensibly decreased. See Figure 1. This also happens when leaving unplugged one coil.

Investigations to be continued. So we leave all the four WE coils unplugged for tests

GPS Coil name Freq (Hz) Amplitude to Coil (Vrms) DARM @ fundamental DARM @ II Harmonic NOTE
1225622466           quiet
1225622616 NE, DR 15 0.884 1E-5 1E-7  
1225622813 NE, DR 15 x 0.5 0.5E-5 0.2E-7  
1225623242 NE, UR 15 0.871 7E-5 1E-7  
1225623559 NE, DL 15 same 7E-7 1E-7  
1225623811  NE, UL 15 same 4E-6 1E-7  
1225624030 NE, DL 15 same 8E-7 1E-7  
1225624151 NE, DL 15 x 2 1.7E-6 4.5E-7  
1225624328  NE, DL 15 x 4 3.5E-6 1.7e-6  
1225631639 NE, DR 15 0.877 1E-5 8E-8 offset in cage position
1225632568 NE, UR 15 same 2 E-5 7 E-8 offset in cage position
1225634751  NE, UR 15 same 2.4E-5 6.6E-8 same
1225634865 NE, DL 15 same 1E-8 7E-8 same
1225635095 NE, UL 15 same 1.7E-6 5.4E-8 same
1225635390 NE, UR 15 same 2E-5 7E-8 same
1225635718 NE, UR 27 same 2.4E-5 2.7E-8 same
1225637892           quiet
1225639862  WE, DR 15 same 3E-6 7E-8  
1225640033 WE, DL 15 same 5.8E-5 7E-8  
1225640446 WE, UR 15 same 3E-5 7E-8  
1225640859 WE, UL 15 same 2.4E-5 7E-8  
1225640973  WE         back to 4 coils connected
1225641075 WE         DL floating
1225641183 WE         short-circuit DL coil, floating --> UNLOCKS
Images attached to this report
Non-image files attached to this report
Comments to this report:
mwas - 20:00 Wednesday 07 November 2018 (43455) Print this report
Figure 1, the improvement with WE coil unplugged seems real. Both h(t) and the DARM spectrum recalibrated by the noise budget agree with each other and are almost a factor two below the level of 1/f^2.5 noise that has been present since July.
Images attached to this comment
vicere - 8:59 Thursday 08 November 2018 (43463) Print this report
ruggi - 15:56 Wednesday 21 November 2018 (43679) Print this report

The effect on DARM has been converted in force, assuming:

- DARMtoH TF@15Hz = 4e-13 (a few percent of difference between NE and WE test has been taken in account, due to the different locking setup)

- HtoF TF@15Hz = 1.12e9

The rms of DARM has been evaluated, then converted in force, then divided by the rms of the applied voltage (0.877).

gps duration TW COIL F/V (nN/V)
1225631939 200 NE DR 2.6
1225634751 100 NE UR 5.1
1225634915 150 NE DL 0.003
1225635145 200 NE UL 0.4
1225639942 70 WE DR 0.6
1225640212 200 WE DL 13.5
1225640562 200 WE UR 6.4
1225640900 70 WE UL 5.4
ruggi - 16:20 Friday 23 November 2018 (43716) Print this report

The results of 15 Hz line injections has been used to project the board common noise (measured by a spectrum analyzer) versus h.

Fig 1: projection of WE noisy p6 onto  hrec; the 1/f^2.3 noise is well explained (20% overestimation)

Fig 2: projection of NE noisy p6 onto hrec with WE unplugged; the contribution is not negligible and a reduction of noise should be visible fixing the problem of common noise. The experiment reported in 43592 did not confirm that, but there was not the possibility to do it in the best way. We will see what happens with the modified p6.

Images attached to this comment
Search Help